1/5x + 3/2 = 3/10
Would I find the common denominator for 1/5x and 3/2, which should be 10? Then do the recipocal of 3/2 to 2/3 and then subtract 3/10-2/3?

Answer by checkley77(12844)

1/5x + 3/2 = 3/10
x/5+3/2=3/10
x/5=3/10-3/2
x/5=(3-3*5)/10
x/5=(3-15)/10
x/5=-12/10
10x=-12*5
10x=-60
x=-60/10
x=-6 ans.
Proof:
-6/5+3/2=3/10
(-6*2+3*5)/(5*2)=3/10
(-12+15)/10=3/10
3/10=3/10